### v1

# Diversity-Enhanced Reasoning for Subjective Questions

Abstract
Large reasoning models (LRM) with long chain-of-thought (CoT) capabilities have shown strong performance on objective tasks, such as math reasoning and coding. However, their effectiveness on subjective questions that may have different responses from different perspectives is still limited by a tendency towards homogeneous reasoning, introduced by the reliance on a single ground truth in supervised fine-tuning and verifiable reward in reinforcement learning. Motivated by the finding that increasing role perspectives consistently improves performance, we propose MultiRole-R1, a diversity-enhanced framework with multiple role perspectives, to improve the accuracy and diversity in subjective reasoning tasks. MultiRole-R1 features an unsupervised data construction pipeline that generates reasoning chains that incorporate diverse role perspectives. We further employ reinforcement learning via Group Relative Policy Optimization (GRPO) with reward shaping, by taking diversity as a reward signal in addition to the verifiable reward. With specially designed reward functions, we successfully promote perspective diversity and lexical diversity, uncovering a positive relation between reasoning diversity and accuracy. Our experiment on six benchmarks demonstrates MultiRole-R1’s effectiveness and generalizability in enhancing both subjective and objective reasoning, showcasing the potential of diversity-enhanced training in LRMs. [^1] The code and data of this work is available at https://github.com/toward-agi/diverse-o1.git.
Diversity-Enhanced Reasoning for Subjective Questions
Yumeng Wangthanks: Equal contribution.    Zhiyuan Fanthanks: Project leader of student team.    Jiayu Liu     Yi R. (May) Fungthanks: Corresponding author Hong Kong University of Science and Technology ywanglu@connect.ust.hk       yrfung@cse.ust.hk
1 Introduction
Advances in OpenAI o1 and DeepSeek R1 style models Jaech et al. (2024); DeepSeek-AI et al. (2025) with long reasoning chains Wei et al. (2023) have significantly enhanced the models’ capability in various reasoning tasks Yu et al. (2025); Wu et al. (2024b). Extended reasoning chains enable models to systematically divide and conquer complex problems, iteratively refining intermediate results, and methodically verifying the outputs. The DeepSeek R1 style models reached performance comparable to human experts in many objective reasoning tasks such as math Cobbe et al. (2021); Wang et al. (2025a); Guo et al. (2025) and coding Jain et al. (2024). This success highlights that scaling up the test time computation can significantly boost performance on complex, structured tasks relative to standard single-step prompting Jaech et al. (2024); Wei et al. (2023).
Figure 1: A subjective reasoning task example which demonstrates that the perpective shifts may lead to different answers. We illustrate that diverse perspective thinking benefits subjective question answering.
Beyond objective tasks with a single correct answer, there is a growing interest in applying Large Reasoning Models (LRMs) to subjective or open-ended problems Sarıtaş et al. (2025); Guo et al. (2024). This homogenization stems from standard training pipelines which reduce subjective problems to a single ground truth for both supervised learning and reinforcement learning, thereby penalizing diverse perspectives. As illustrated in Figure 1, due to the single inherent thinking process of LRM, the model tends to rely on homogeneous reasoning trajectories without exploring alternative perspectives, hindering their capacity to incorporate diverse reasoning strategies. Prior works’ solutions mainly fall into two categories: (1) Research on multi-agent debates has demonstrated that incorporating diverse role perspectives substantially improves performance on subjective tasks Aoyagui et al. (2025); Cheng et al. (2024); Liu et al. (2025b), but these methods require multiple interacting models; (2) some approaches ensemble solutions from varied decoding paths Wang et al. (2023), yet the stochastic nature of these paths may not necessarily introduce a broader range of perspectives Naik et al. (2024).
To address this gap, we propose to enhance the diversity of the DeepSeek R1 style reasoning chain, by dynamically integrating diverse role perspectives for subjective question answering, including tasks such as moral dilemmas, opinion-based questions, ambiguous question answering, and so on. To begin with, we conduct a pilot study to confirm two key factors: the effectiveness of long reasoning chains in subjective reasoning tasks, and the impact of varying the number of role perspectives incorporated during reasoning. Our findings show that sequential test-time scaling performance improves as reasoning chains are lengthened, but begins to decline once they surpass an optimal length. Similarly, we identify that there exists an optimal range for the number of role perspectives in the reasoning chain. Building upon these insights, we propose a novel framework MultiRole-R1 that enhances the diversity of long chains of thought. Specifically, we optimize the role perspective diversity and the reasoning diversity.
MultiRole-R1 starts with a novel data construction pipeline combining parallel scaling with sequential scaling to effectively incorporate a broader array of role perspectives into the long reasoning chains. We subsequently refine the model through supervised fine-tuning (SFT), instructing the model to learn the multi-role reasoning format and enhance perspective diversity. Furthermore, to enhance the diversity of the reasoning process, we introduce a carefully designed diversity reward function implemented through a multi-role Generalized Reward Policy Optimization (GRPO) framework, taking diversity as an additional reward signal. By training on subjective questions only, MultiRole-R1 boosts the performance of both subjective and objective reasoning tasks. Our contribution can be summarized as follows:
  • We perform a pilot analysis that reveals the scaling law of reasoning length and role perspectives in subjective reasoning.
  • We propose MultiRole-R1, a training paradigm that incorporates unsupervised SFT data generation and GRPO with reward shaping, which effectively promotes perspective and lexical diversity. By training on subjective questions solely, MultiRole-R1 achieves state-of-the-art performance in both subjective and objective reasoning benchmarks.
  • Our experiments and analysis highlight an interesting phenomenon: optimizing for diversity often aligns with, and can even enhance, the accuracy objective. This synergistic relationship underscores our method’s effectiveness and its robust generalizability to other domains.
2 Related Work
Test-Time Scaling
Recent advancements, such as the success of Deepseek-R1 DeepSeek-AI et al. (2025) in demonstrating robust long-chain-of-thought (CoT) reasoning, have brought significant attention to test-time scaling Wang et al. (2025c); Liu et al. (2025a); Li et al. (2025b). Current approaches to test-time scaling can be broadly classified into four categories Zhang et al. (2025). Sequential scaling, as explored by Madaan et al. (2023) and Xiang et al. (2025), iteratively refines a model’s state and output, with each step building on the previous to form a coherent chain of computations. In contrast, parallel scaling Wu et al. (2025); Wang et al. (2024a) leverage parallelism to expand the coverage of reasoning chain, thereby increasing the likelihood of identifying correct solutions. Additionally, hybrid scaling, as proposed in Wang et al. (2024b) and Zhang et al. (2024), combines parallel generation of diverse options with sequential evaluation and refinement to enhance reasoning capabilities. Finally, internal scaling Muennighoff et al. (2025); Ye et al. (2025) empowers models to autonomously allocate computational resources during inference, moving away from reliance on externally guided strategies. While these approaches have demonstrated impressive performance improvements, they primarily focus on enhancing models’ reasoning capabilities in mathematical or coding tasks which contain fixed and verifiable answers. Such methods largely overlook the importance of promoting diversity in reasoning for open-ended questions, which require consideration from multiple perspectives.
Subjective Tasks and LLM Role-Playing
Subjective task is a type of task that differs from objective tasks, such as commonsense reasoning, code generation and arithmetic reasoning Wang et al. (2025b) -which involve questions with definitive correct or incorrect answers. Subjective task, due to its open-ended nature, has answers that may change with perspective-shifts or context change Wang et al. (2025b); Jentzsch and Kersting (2023); Wu et al. (2024a). This includes culture-related question answering Huang and Yang (2023); DURMUS et al. (2024); Huang et al. (2025b), subjective language interpretation Jones et al. (2025), ethical question answering Hendrycks et al. (2021), creative question answering Lu et al. (2024), and so on. LLM role-playing Shao et al. (2023), including multi-role discussion Wang et al. (2024c); Du et al. (2023); Liang et al. (2024) is one of the most used techniques to study subjective task. It features specialized AI systems to simulate assigned personas, ranging from real-life figures to fictional characters Chen et al. (2024); Li et al. (2025a). The capacity to simulate different personas has been shown to elicit human-like responses, and introduce more diverse and creative reasoning paths Naik et al. (2024) when solving LLM-related tasks Wang et al. (2024d). In this work, we adopt a parallel multi-role reasoning at test time to enable the model think from a diverse perspective via independent reasoning paths.
3 Pilot Analysis
Figure 2: (a) The performance of Deepseek-r1-distill-qwen-7b DeepSeek-AI et al. (2025) under different reasoning length settings across different datasets. The bar chart shows that longer reasoning chains result in higher accuracy on subjective tasks.  (b) The performance of Deepseek-r1-distill-qwen-7b DeepSeek-AI et al. (2025) under different reasoning length settings across different datasets.  (c) Demonstration of the number of distinct opinions increases as more roles are involved in a single reasoning chain.
Drawing on evidence that diverse perspectives enhance performance on subjective tasks Muscato et al. (2025); Hayati et al. (2024), we introduce perspective diversity into Deepseek R1-style reasoning chains via test-time scaling Snell et al. (2024). We employ two complementary strategies: sequential scaling to deepen reasoning through iterative reflection Zeng et al. (2025), and parallel scaling to enhance exploration by generating multiple reasoning paths Liu et al. (2025c); Rodionov et al. (2025). We evaluate this approach on three subjective reasoning datasets: GlobalOpinion QA DURMUS et al. (2024), BBQ Parrish et al. (2022), and ETHICS Hendrycks et al. (2021) across four decoding strategies adapted from Jiang et al. (2025), with examples provided in Appendix 7:
  • Zero think: Force the model to respond without thinking, i.e. “<think></think>”.
  • Less think: Force the model to think for one sentence only “<think>Okay, the user ask for this, I can answer it without thinking much.</think>”
  • Regular think: Let the model start with “<think>” and ends its thinking naturally.
  • More think: Starts with a regular think. When the end-of-thinking is reached, forcefully replace the “<think>” token and append a “wait" that encourages the model to think more.
As shown in Figure 2 (a), increasing the amount of test-time reasoning through scaling indeed yields performance gains. We thus incorporate More think as our baseline.
Scaling Law of Reasoning Length
Having established the effectiveness of test-time scaling, a natural follow-up question is whether longer reasoning chains always result in better performance. However, prior studies on mathematical and commonsense reasoning have shown that excessively long reasoning chains do not necessarily lead to improved outcomes Sui et al. (2025); Ballon et al. (2025); Qin et al. (2025). Therefore, to control the reasoning length for sequential scaling leveraging concepts from budget forcing Muennighoff et al. (2025); Huang et al. (2025a), we conduct an investigation experiment aimed at roughly estimating the optimal reasoning length specifically for subjective problems. We adhere to the experimental setup proposed by Jiang et al. (2025) and experiment on the optimal number of replacing the end-of-thinking delimiter. As shown in Figure 2 (b), increasing the number of "Wait" tokens generally leads to performance improvements across most tasks. For GLOQA, performance continues to improve with more “Wait” insertions, whereas for BBQ and ETHICS, the gains peak around three insertions and diminish or even degrade beyond that point. Based on this observation, we adopt three “Wait” tokens as a balanced setting that achieves clear performance gains without excessive computational overhead. We refer to this configuration as the More think strategy, which not only fosters more cautious and reflective reasoning, but also enables longer reasoning chains capable of supporting more nuanced and diverse role-based perspectives.
Scaling Law of Role Perspectives
We explore the relationship between the number of roles and the answer diversity and consensus accuracy on subjective QA tasks. Intuitively, each additional role contributes a new viewpoint, expanding the solution space. Lu et al. (2024) shows multi-agent debates with 3 fixed roles achieve optimal creativity, but scaling different role perspectives in a single reasoning chain at test time remains unexplored. We conduct a pilot analysis that directly instructs the model to think from $n$ different viewpoints, with $n$ from 1 to 6. Results in Figure 2 (c) show that the number of distinct opinions increases as more roles are involved. In general, we observed that the number of distinct opinions plateaus as $n=4$. Hence, in the following study, we control the number of generated roles $n=2,3,4$ for efficient generation.
Figure 3: Illustration of our MultiRole-R1framework. Stage 1: we collect our data by utilizing budget forcing (sequential scaling) and contrastive role reasoning, sampling N outputs from the decoder (parallel scaling). Stage 2: we automatically construct our training data that integrates various perspectives into a single reasoning chain, and then fine-tune the model to follow the multi-role reasoning format. Stage 3: we utilize GRPO with reward shaping to optimize the reasoning accuracy and diversity. Two accuracy evaluation approaches are applied, depending on whether the ground-truth varies by roles. By taking diversity as an additional reward, we encourage exploration of the reasoning chain, expanding the answer search space.
4 Method
As illustrated in Figure 3, our framework consists of three stages: parallel multi-role reasoning, multi-role finetuning and multi-role reinforcement learning. Formally, given an input question $\mathcal{Q}$ and a reasoning model $\mathcal{M}$, our goal is to diversify the reasoning path $\mathcal{T}$.
4.1 Parallel Multi-Role Reasoning
Multi-Role Exploration
To model multi-perspective reasoning, we first identify a set of $n$ context relevant roles (e.g., domain experts, stakeholders, or personas) through few-shot prompting Brown et al. (2020), denoted as =$\mathcal{R}=\{R_{1},R_{2},...,R_{n}\}$. In particular, we prompt the model to generate roles that may have conflicting viewpoints. The motivation for this is to explore diverse available perspectives. Given candidate roles $\mathcal{R}$, we define the selection probabilities:
$\begin{split}P(R_{i}|\mathcal{Q})=softmax(\mathbb{E}[\mathcal{M}(R_{i}|\mathcal{Q})]\\ +\lambda\mathbb{E}_{R_{i}}[1-\text{sim}(R_{i},R_{j})]),\end{split}$ (1)
where $sim(R_{i},R_{j})=cos(\mathbf{h}_{R_{i}},\mathbf{h}_{R_{j}}|\mathcal{Q})$ and $\mathbf{h}$ denotes the embedding of the LLM $\mathcal{M}$. These terms represent that, in the role selection process, we prioritize role answers that are relevant to the question $\mathcal{Q}$ and contrastive to the existing opinions.
Self-Consistency Filtering
For each role $R_{i}$, we generate $k$ candidate reasoning paths $\mathcal{M}(Q,R_{i})=\mathcal{T}_{R_{i}}=\{T_{R_{i}}^{(1)},T_{R_{i}}^{(2)},...,T_{R_{i}}^{(k)}\}$. This is achieved by adjusting the sampling temperature $\tau=1$ to encourage diverse reasoning paths. To ensure the coherence among different responses of each role, we then apply self-consistency filtering Chen et al. (2023); Wang et al. (2025d) through majority voting and only keep the most consistent answer:
$\hat{T}_{R_{i}}=argmax\sum_{j=1,T\in\mathcal{T}_{R_{i}}}^{k}\mathbbm{1}(T\equiv T_{R_{i}}^{(j)}),$ (2)
where $\mathbbm{1}$ is the indicator function and $\equiv$ denotes semantic equivalence (e.g. same roles give different answers). This approach extends ensemble methods by decoupling role-specific reasoning trajectories, ensuring that conflicting viewpoints remain independently generated and self-consistent.
4.2 Multi-Role Finetuning
To enhance our model’s reasoning capabilities in both accuracy and diversity, we first fine-tune the base model with self-consistency filtered data. This initial stage teaches the model to adopt a multi-role reasoning format, which involves shifting perspectives throughout its reasoning process. To further promote this diversity, our supervised fine-tuning (SFT) dataset includes 968 examples specifically designed for varied role generation.
Sequential Multi-Role Merging
Given $m$ filtered role perspectives $\{\hat{T}_{R_{1}},...,\hat{T}_{R_{m}}\}$, we generate random combination of role orderings $\Pi$ to avoid the effect of position bias Zheng et al. (2023a). For example, given a multi-role combination $\pi=\{R_{i},R_{j},R_{k}\}$, we construct the training data as follows:
$\mathcal{D}_{\text{train}}=\bigcup_{\pi\in\Pi}\{(\mathcal{Q}\oplus\hat{T}_{R_{i}}\oplus\hat{T}_{R_{j}}\oplus\hat{T}_{R_{k}})\mid\pi\}.$ (3)
We consider two merging strategies depending on task type to allow dynamic integration of role reasoning traces while respecting the nature of the reasoning task: (1) Divergent merging: when the roles are supposed to give different answers, we encourage diverse perspectives to maximize coverage of the searching space. Final prediction is derived via weighted aggregation over differing viewpoints. (2) Convergent merging: when the roles are supposed to give consistent answers, they reach consensus via majority voting within a single sequence, which is in nature soft majority voting within a single sequence.
Multi-Role Supervised Finetuning
To ensure the quality of the SFT data, we apply several filtering strategies to the merged dataset. To mitigate verbosity bias Zheng et al. (2023b) and reasoning shortcut behaviors, we remove entries in the top and bottom 10th percentiles of answer length. We also discard instances with formatting errors or invalid string patterns. This results in the final 2,700 SFT training data. Additionally, since the gold labels are accessible in the original dataset, we apply a supervised, ground-truth-guided filtering approach as a comparison to our unsupervised self-consistency method. Results of this comparison are presented in Section 7.1.
4.3 Multi-Role Reinforcement Learning
We adopt Group Relative Policy Optimization (GRPO) Shao et al. (2024) for Multi-Role reinforcement learning, which is trained on top of the SFT model. GRPO optimizes the policy by sampling a group of candidate outputs for each prompt and comparing their reward. We incorporate two types of rewards: an accuracy reward $R_{\text{acc}}$ provided by a verifiable reward model that checks answer correctness, and a diversity reward $R_{\text{div}}$ computed from the input text as a shaping signal. The total shaped reward is given by $R=0.9*R_{\text{acc}}+0.1*R_{\text{div}}$. This follows the reward-shaping paradigm Ng et al. (1999), where the auxiliary $R_{\text{div}}$ guides learning without changing the optimal policy.
During training, we observe a synergetic effect of optimizing the diversity and accuracy objectives. This also mitigates issues observed in the SFT baseline, such as excessive verbosity and repetitive reasoning Toshniwal et al. (2025). Finally, note that GRPO computes group advantages $A_{1},A_{2},\dots,A_{G}$ instead of standard reward, which is given by:
$\hat{A}_{i,t}=(R_{i,t}-\mu)/\sigma,t\in\{1,\dots,|G|\}.$
Hence, a group with uniform rewards (all 0s or all 1s) would give zero advantage and stall learning. By adding the diversity term, we ensure intra-group reward variance, enabling informative gradients and continued optimization.
5 Experiment
Model BBQ GLOQA CALI (O.O.D) ETHICS CSQA (O.O.D) GSM8K (O.O.D)
Acc. Div. Acc. Div. Acc. Div. Acc. Div. Acc. Div. Acc. Div.
(R1-Distill-Qwen-7B)
Zero-shot CoT 62.45 56.02 32.62 65.88 50.30 52.22 51.82 36.14 63.06 83.83 80.48 68.08
Self-Refine 74.08 73.13 43.13 59.88 50.76 66.09 52.19 37.36 54.02 77.61 87.01 80.37
Role-Play 73.61 74.68 41.67 77.75 52.69 67.43 50.83 37.89 55.07 76.20 85.66 72.87
More think 80.76 80.44 36.42 86.90 60.45 78.82 64.44 81.53 64.50 85.85 82.05 81.79
Ours SelfConsis SFT 85.88 81.67 43.13 85.58 67.35 78.94 67.45 82.19 66.88 83.10 80.62 74.87
Ours SelfConsis SFT+DPO 86.41 60.43 44.20 61.17 68.19 64.09 67.28 68.51 67.24 69.83 81.51 67.56
Ours SelfConsis SFT+GRPO 94.30 85.52 47.22 87.46 70.83 82.15 69.50 85.40 69.43 86.85 85.58 82.16
MultiRole-R1 SelfConsis SFT+GRPO(RS) 94.50 86.25 49.10 89.67 70.85 83.31 66.83 87.27 66.94 87.96 87.36 82.46
(R1-Distill-Llama-8B)
Zero-shot CoT 80.89 79.92 38.41 87.07 60.84 73.98 62.46 79.44 67.21 83.39 78.87 76.52
Self-Refine 74.20 75.85 43.19 81.11 61.95 78.87 60.96 80.17 63.77 82.61 80.95 81.24
Role-Play 74.40 80.91 44.87 83.02 62.70 77.32 64.24 79.78 67.32 82.27 77.33 75.02
More think 88.20 84.11 44.04 87.19 64.41 80.30 68.06 83.99 70.42 84.73 83.30 84.12
Ours SelfConsis SFT 89.69 82.64 48.17 87.26 70.05 79.77 70.56 81.36 70.86 83.88 86.02 81.53
Ours SelfConsis SFT+DPO 90.52 80.43 48.89 80.97 69.81 76.62 71.22 83.64 71.28 82.71 86.34 81.81
Ours SelfConsis SFT+GRPO 94.47 85.75 48.55 89.36 69.26 83.37 75.63 87.89 73.71 87.96 87.49 85.31
MultiRole-R1 SelfConsis SFT+GRPO(RS) 95.55 89.58 49.06 91.78 71.48 90.55 75.84 96.54 75.12 92.98 89.79 88.45
(R1-Distill-Qwen-14B)
Zero-shot CoT 85.01 68.06 36.82 79.18 75.05 71.83 73.63 72.20 75.85 83.09 85.58 70.68
Self-Refine 90.42 80.13 49.04 69.40 71.28 78.22 76.48 83.22 76.55 82.31 84.73 80.24
Role-Play 91.18 81.87 49.90 75.73 67.41 70.74 77.16 75.30 75.71 79.05 91.50 76.60
More think 94.57 80.67 41.60 84.04 75.90 76.81 79.36 83.33 79.36 81.77 88.76 80.94
Ours SelfConsis SFT 94.40 75.06 50.98 81.04 76.08 73.65 81.45 71.34 81.50 77.60 91.61 91.62
Ours SelfConsis SFT+DPO 94.98 67.21 51.33 65.03 75.82 71.71 81.88 42.32 79.82 69.41 90.92 68.69
Ours SelfConsis SFT+GRPO 95.98 86.88 51.73 90.33 75.65 84.92 83.50 89.42 81.19 89.64 91.87 86.36
MultiRole-R1 SelfConsis SFT+GRPO(RS) 97.50 90.17 53.98 91.32 76.50 89.08 86.00 92.89 82.00 91.61 93.43 87.24
(Qwen3-8B)
Zero-shot CoT 91.71 70.10 42.13 60.99 73.40 57.43 72.29 76.68 80.81 57.84 85.41 70.49
Self-Refine 88.93 53.07 45.25 85.00 69.40 48.16 70.64 49.60 69.22 50.99 84.91 82.31
Role-Play 89.77 41.58 47.57 54.89 70.26 50.49 70.67 47.83 72.98 48.17 93.58 81.93
More think 95.18 74.20 43.39 78.98 75.10 68.44 78.26 72.45 81.20 73.07 90.02 75.07
Ours SelfConsis SFT 94.05 74.02 50.32 77.07 75.96 72.78 78.39 68.35 81.00 73.50 91.62 70.23
Ours SelfConsis SFT+DPO 94.21 74.83 50.68 74.14 76.10 69.58 78.09 50.31 80.45 64.87 91.84 65.32
Ours SelfConsis SFT+GRPO 95.91 85.47 51.37 87.74 77.83 80.36 79.82 86.84 81.19 86.40 91.97 85.98
MultiRole-R1 SelfConsis SFT+GRPO(RS) 96.98 88.15 51.72 89.88 77.95 83.84 81.95 89.09 82.10 87.93 94.98 86.93
Table 1: Main results of the baselines (specified in Section 5) and our proposed method. Acc. is the accuracy of the task (in %) and Div. measures the length normalized diversity score of the reasoning chain (in %). We include two ablations of MultiRole-R1, including SFT on self-consistency filtered data only (Ours SelfConsis SFT), and also SFT with vanilla GRPO (Ours SelfConsis SFT + GRPO). “GRPO(RS)” represents GRPO with reward shaping, which is used in MultiRole-R1. O.O.D. denotes the datasets that are for testing only. Detailed decomposition of the diversity score is shown in Appendix E.
5.1 Datasets
We train our model on three subjective tasks: ambiguous question answering (BBQ by Parrish et al. (2022)), opinion-based QA (GlobalOpinionQA by DURMUS et al. (2024)), and ethical dilemma (ETHICS by Hendrycks et al. (2021)). To evaluate the effectiveness and generalizability of our approach, we test on three additional datasets: cultural natural language inference (CALI by Huang and Yang (2023)), commonsense reasoning (CSQA by Talmor et al. (2019)), and mathematical reasoning (GSM8K by Cobbe et al. (2021)). Specifically, for the testing data, CALI consists of the out-of-domain (O.O.D.) subjective questions, while CSQA and GSM8K consist of O.O.D. objective questions. It is worth noting that among these benchmarks, GLOQA and CALI have role-dependent ground truths, whereas the others rely on a single ground truth for all roles.
5.2 Baselines
In-Context Learning
We first incorporate the following in-context learning Brown et al. (2020) settings: (1) Zero-Shot CoT Kojima et al. (2023); Wei et al. (2023), (2) Role Playing Prompting Kong et al. (2024) and (3) Self-Refine Prompting Madaan et al. (2023).
More Think
As observed by Muennighoff et al. (2025), extending the reasoning chain length can further enhance the reasoning capabilities of o1-style models. In MultiRole-R1, this is achieved by suppressing the end-of-thinking token and appending a continuation string (e.g., “wait, I need to think from {role}’s perspective") to encourage extended reasoning from a different role perspective. In the more think baseline, we employ a reasoning length three times longer than regular think, as it offers a balance between efficiency and accuracy based on our pilot analysis.
Supervised Finetuning
We perform supervised finetuning on the base model on the self-consistency filtered dataset of size 2,700. Our SFT training and evaluation are conducted via Llama-Factory Zheng et al. (2024).
Direct Preference Optimization
We introduce another SFT+RL pipeline as an ablation to MultiRole-R1. In this setting, DPO is applied to the self-consistent SFT model, using ground-truth-hinted role answers as positive samples and inconsistent role answers as negative samples.
5.3 Models
Our experiment is performed on a range of open-source LRMs including R1-Distill-Qwen-7B, R1-Distill-Llama-8B and R1-Distill-Qwen-14B models DeepSeek-AI et al. (2025), and Qwen3-8B Qwen Team (2025) with reasoning mode.
Figure 4: (a) We plot the accuracy against the diversity for each setting (e.g. zero-shot, role-play, morethink, etc.) for each dataset. We observe a positive linear correlation between the two metrics.  (b) Qualitative example of the 34 most frequent roles in the training data generated by LRMs. A more detailed visualization is presented in Figure 5.
5.4 Metrics
Accuracy
Taking into account the subjective nature of role-based reasoning where the ground truth for subjective questions may vary across different roles, we adopt two different perspective merging strategies during evaluation.
(1) Divergent Merging: for tasks such as CALI and GLOQA, each role $i$’s answer $a_{i}$ is compared with the corresponding ground truth $g_{i}$, where the divergent accuracy is given by:
$Acc_{div}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[a_{i}=g_{i}].$
(2) Convergent Merging: datasets like BBQ, ETHICS, CSQA and GSM8K have answers invariant with role-perspectives. We aggregate different role’s answer to obtain a consensus, and then compare it to the ground truth:
$\hat{a}=argmax\sum_{i}\mathbbm{1}(a_{i}=\hat{a}),Acc_{con}=1[\hat{a}=g].$
Diversity
To quantify the diversity of model-generated reasoning, we design a composite metric that captures linguistic variation across multiple dimensions. Inspired by prior work on lexical and entropy-based diversity in natural language generation Li et al. (2016); Tanaka-Ishii and Aihara (2015), our metric is a weighted sum of eight complementary diversity signals, including lexical, token entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, distinct N-gram and function word diversity. Formal definition of the diversity metrics can be found in Appendix E. Formally, we express the final combined diversity score in the formula below:
$D_{final}=0.15\sum D_{i}+0.1\sum D_{j},$ (4)
where: $D_{i}\in\{D_{lex},D_{ent},D_{pat},D_{bi}\}$ and $D_{j}\in\{D_{len},D_{adj},D_{yule},D_{func}\}$. The choice of the weighting is further illustrated in Section 7.2.
6 Results and Analysis
6.1 Main Results
Table 1 shows that with multi-role SFT and GRPO with reward shaping, MultiRole-R1 achieves superior performance in both subjective and objective reasoning tasks, surpassing More Think by an average of 7.6% and 3.8% in accuracy and diversity. Notably, by training solely on subjective tasks, we enhance both the subjective and objective reasoning abilities of models.
For the choice of RL algorithm, we found that GRPO yields more accurate and diverse responses than DPO. We attribute this to a fundamental mismatch between DPO’s training format and the nature of our task. Subjective questions do not have single "correct" answers, yet DPO requires a rigid "preferred-rejected" pair for training. These results therefore demonstrate the superiority of GRPO for this context and validate the effectiveness of explicitly optimizing for diversity.
Interestingly, Table 9 to 12 in the appendix show that our approach results in a decrease in answer tokens compared to SFT, suggesting that mere verbosity does not lead to more accurate answers.
6.2 Domain Generalization
As shown in Table 1, besides in-domain subjective questions that are included in the training set, MultiRole-R1 successfully boosts the performance in all the benchmarks. This shows our approach’s generalizability is two-fold: generalizability to other subjective questions, and also generalizability to objective questions. We reveal that by only training on three subjective datasets, the model gains a generalized understanding of role generation and role reasoning, and has a near 10% accuracy gain in subjective tasks like cultural natural language inference in CALI. Moreover, without any training in objective questions, MultiRole-R1 significantly enhances the models’ performance on commonsense reasoning and math reasoning benchmarks. This demonstrates that subjective question training can potentially enhance the performance of objective questions. An explanation is that the nature of subjective questions pushes the model to gain insights from various perspectives, which collectively guides the model to approach the optimal solution.
Model BBQ GLOQA CALI ETHICS CSQA GSM8K
R1-Distill-Qwen-7B 94.2 45.5 89.2 98.6 92.4 58.4
R1-Distill-Llama-8B 89.0 48.0 82.4 83.6 85.6 64.9
R1-Distill-Qwen-14B 82.0 59.0 60.5 77.2 57.4 79.2
Qwen3-8B 76.0 65.4 89.3 73.0 82.5 33.4
Table 2: The task break-down correlation coefficient (in %) between accuracy and diversity.
6.3 Accuracy-Diversity Correlations
Our analysis reveals a positive correlation between accuracy and diversity (Figure 4a). This relationship is reinforced by a strong per-task correlation between diversity and accuracy (0.9), which significantly exceeds the correlation with response length (Table 2). These results suggest that performance gains in modern reasoning models like o1 or R1 are driven by a scaling law of diversity, rather than superficial verbosity. Crucially, this demonstrates that optimizing for diversity does not compromise correctness but rather enhances accuracy. One possible explanation is that optimizing for diversity can serve as a useful inductive bias, enabling the model to explore a broader solution space and discover more accurate, perspective-aligned answers in subjective tasks.
Model BBQ GLOQA CALI (O.O.D) ETHICS CSQA (O.O.D) GSM8K (O.O.D) Alignment
Human Div. Human Div. Human Div. Human Div. Human Div. Human Div.
R1-Distill-Qwen-7B 8.95 86.25 9.62 89.67 7.88 83.31 8.24 87.27 8.51 87.96 7.13 82.46 0.88
R1-Distill-Llama-8B 7.94 89.58 8.33 91.78 8.78 90.55 9.82 96.54 9.15 92.98 7.21 88.45 0.93
R1-Distill-Qwen-14B 8.81 90.17 9.52 91.32 8.43 89.08 9.65 92.89 9.18 91.61 8.24 87.24 0.95
Qwen3-8B 8.51 88.15 9.62 89.88 7.42 83.84 9.25 89.09 8.23 87.93 8.88 86.93 0.89
Table 3: We use human ratings as a reference to set the weights. The alignment score of human rating and the combined diversity score shows that our metric highly aligns with human preference.
6.4 Role Coverage
We leverage the LLM itself to generate roles pertinent to the context of the question. In total, there are 968 distinct roles in the training data, enhancing the perspective diversity of the LRM. We leverage the LLM itself to generate roles relevant to the context of each question. These roles demonstrate broad coverage, including different moral philosophies, nationalities, identity groups, and specific individuals pertinent to the questions. Figure 4 (b) shows a qualitative example of the most frequently occurred roles.
7 Discussion
7.1 Filtering Strategy of the Training Data
Our methodology compares two data sampling strategies for Supervised Fine-Tuning (SFT): an unsupervised self-consistency approach and a supervised method that leverages ground-truth agreement. Table 4 reports the test performance on equal-sized datasets generated by each strategy. The results indicate that while self-consistency filtering can yield slightly lower accuracy in some cases, its performance is broadly comparable to, and can even outperform, supervised filtering. This highlights that using noisy, unsupervised data is sufficient and viable for our tasks Wang et al. (2025d).
BBQ GLOQA CALI ETHICS CSQA GSM8K
(R1-Distill-Qwen-7B)
Consis. Filter 85.55 47.13 67.35 67.45 66.88 80.62
GT Filter 88.40 45.55 65.95 68.44 68.45 80.63
(R1-Distill-Llama-8B)
Consis. Filter 89.69 48.17 72.05 70.56 70.86 83.30
GT Filter 87.44 49.29 69.27 72.15 71.06 84.20
(R1-Distill-Qwen-14B)
Consis. Filter 94.40 50.98 76.98 81.45 80.50 91.61
GT Filter 94.88 52.29 76.28 81.57 80.79 91.28
(Qwen3-8B)
Consis. Filter 94.05 50.32 75.96 78.39 81.00 91.62
GT Filter 94.80 51.07 76.15 80.19 82.13 91.85
Table 4: The evaluation accuracy after finetuning on (1) consistency filtering and (2) ground-truth hinted sampling data.
7.2 Design of the Diversity Weighting
In equation (4), we design our diversity metric as a combined score of lexical diversity, token entropy diversity, etc. Our guiding principle of diversity metric design is to balance the weights of different diversity factors and make it align with human intuition as much as possible. We present Table 3 showing qualitative alignment scores between the metrics and human diversity ratings. We manually label 60 outputs from different settings, from a diversity scale of 1 to 10. Taking the human rating as a reference, we set the weight of the diversity metrics such that the correlation coefficient between the overall diversity and the human rating is highly aligned. Table 3 shows that our diversity metric has a high alignment score with human ratings, showcasing the reasonableness of our weighting.
7.3 Advanced Math Reasoning
To further validate the robustness of our pipeline on unseen objective tasks, we introduced the AIME 2024 benchmark, which is considerably more challenging than GSM8K. Notably, despite being trained exclusively on subjective datasets, our method demonstrates a consistent O.O.D. performance improvement on AIME, similar to the gains observed on GSM8K.
Model r1-dist-qwen7B r1-dist-llama8B r1-dist-qwen14B qwen3-8B
Zero-Shot 55.5 50.4 69.7 76.0
Self-Refine 55.9 50.8 70.6 77.5
Role-Play 56.3 53.4 69.1 76.6
More think 58.6 54.0 71.4 78.8
SFT 58.9 56.8 70.2 78.3
SFT+GRPO 62.7 56.7 72.8 79.6
MultiRole-R1 63.2 58.1 73.3 80.1
Table 5: Accuracy (in %) on AIME 2024 benchmark, where MultiRole-R1 consistently outperforms other baselines.
8 Conclusion
We introduce MultiRole-R1, a novel framework that enhances the reasoning capabilities of Large Reasoning Models (LRM) on subjective tasks. Our framework enables models to think from multiple perspectives before providing a final answer, enabling diverse answers for subjective questions. To achieve this, we employ Group Relative Policy Optimization (GRPO) with a shaped reward that combines a verifiable accuracy signal with a diversity metric. Furthermore, through training on subjective questions solely, we are able to enhance the models’ performance on both subjective and objective reasoning. Our analysis reveals that diversity is a more consistent indicator of performance than response length, showing a strong per-task correlation with accuracy that significantly outweighs the correlation with length. This suggests that the improvements in modern reasoning models like o1 or R1 are driven by a scaling law of diversity, not just superficial verbosity. Our work underscores the importance of diversity at both perspective and lexical levels, paving new directions for future research on diversity-enhance training of LRMs.
Limitations
One of the main limitations of our study is that due to the constraints of computational resources, we are unable to scale to larger models, e.g. models of 70B. For the multirole supervised finetuning process, since the filtering results after self-consistency are still noisy, the final alignment performance may be constrained by the noisy labels in the positive samples. Our role generation solely depends on the model itself, which may not accurately represent the opinion of the real social community. Future work can further investigate the model’s culture and opinion alignment to different demographics, by incorporating the value of a broader demographic of users. In addition, beyond text-only data, extending diversity-enhanced reasoning in the multimodal domain through thinking with images Su et al. (2025) stands as another underexplored yet meaningful research area to investigate further.
Ethics Statements
Our study focuses on subjective question answering, which encourages LLM’s answers to include more diverse social identities, moral values, and nationalities. To achieve multiple-perspective reasoning, our approach utilizes the model’s own outputs to perform role generation. We also unleash the LLM’s role-playing ability to generate diverse opinions based on the role perspectives. By reducing dependence on manual labeling, this method enhances the fairness, scalability, and inclusivity of AI, which will further the democratization of LLMs across broader demographics.
Author Contribution Statements
Yumeng Wang (YW), Zhiyuan Fan (ZF), and Jiayu Liu (JL) jointly completed the project and contributed equally. In particular:
  • ZF proposed the idea of perspective and lexical diversity for subjective questions.
  • ZF, YW, and JL conceived of the idea of data construction presented in this paper.
  • YW and JL developed and completed the current pipeline of the proposed method.
  • YW and JL performed the experiments in the pilot analysis, supervised fine-tuning, and ZF optimized the inference speed. ZF performed the GRPO training and designed the reward shaping.
  • YW designed the figures and tables. All three co-first authors contributed to writing the paper, with YW leading the writing effort and discussing with ZF and JL.
Yi R. (May) Fung advised the project, participated in the discussion, provided detailed and important suggestions, and helped proofread the entire paper.
Appendix A Training and Testing Data Split
We report the number of merged data constructed, and the number of data remaining after applying the filtering strategy in Table 6. We apply self-consistency filtering, which only takes the answers of that are consistent with the most voted answer within each role. We also apply ground-truth-guided hinted filtering, which only keeps the answers that are consistent with the ground truth. To ensure a fair comparison, we ensure that the number of data left after each filtering strategy is the same.
BBQ GLOQA ETHICS CALI CSQA GSM8K
Merged 3883 4659 2400 - - -
+Consis. filter 1000 1200 500 - - -
+GT filter 1000 1200 500 - - -
Test set 831 999 500 500 496 1000
Table 6: Statistic of the number of training data after self-consistency filtering (consistency filter) and ground-truth-guided hinted filtering (GT filter). We also report the number of test data used in the evaluation phase.
Appendix B Qualitative Examples
B.1 Qualitative Examples of Test Time Scaling Set Up
Table 7 shows the different lengths of the reasoning chain in the pilot analysis. Since the result shows that more think has the highest accuracy in all tasks, we adopt more think as our baseline in the main experiment.
Regular think Instruction Please think from diverse perspectives to answer the question.
Respond in the following format:<think>…</think>…
Input < | User | >Is online courses more effective than traditional classroom?
< | Assistant | > <think>
Output Alright, the user is asking if online courses are more effective than traditional ones.
From one perspective, online courses offer flexibility….
</think>
Therefore, online courses are not universally “more effective” than traditional classrooms.
Zero Think Instruction Please think from diverse perspectives to answer the question.
Respond in the following format:<think>…</think>…
Input < | User | >Is online courses more effective than traditional classroom?
< | Assistant | > <think></think>
Output The effectiveness of online courses versus traditional classrooms depends on various factors…
Final Answer: Neither is universally superior
—effectiveness depends on the subject, learner’s style, and institutional support.
Less Think Instruction Please think from diverse perspectives to answer the question.
Respond in the following format:<think>…</think>…
Input < | User | >Is online courses more effective than traditional classroom?
< | Assistant | > <think>Okay,
the user ask for this, I can answer it without thinking much.</think>
Output The “most effective” method depends on individual needs and the learning context…
More Think Instruction Please think from diverse perspectives to answer the question.
Respond in the following format:<think>…</think>…
Input < | User | >Is online courses more effective than traditional classroom?
< | Assistant | > <think>
Output 111 Okay, so the question is whether online courses are more effective than traditional classrooms.
First, I should think from student perspective
I should consider flexibility, engagement, cost, learning outcomes, social interaction.
</think> Wait, but I also need to think from educator’s perspective
Output 222 Let me take teaching effectiveness, student engagement, workload, and feedback quality into consideration.
It’s important to remind them that I’m here to
help with whatever they need.
</think> Wait, but I also need to think from a parent’s perspective
$\cdots$
Output $N$ Employers in tech may value online certifications.
</think>
…Effectiveness ultimately depends on aligning the mode of learning with the goals and context of the stakeholder.
Table 7: The grey yellow, green boxes are the instructions, reasoning chains, and the model response. Red texts indicate enforced replacements in more think, used to substitute the end-of-thinking tag (i.e., </think>).
B.2 Qualitative Example of Different Roles
Figure 5 shows the top 100 most frequent roles in the dataset. The role covers a broad range of groups including but not limited to:
  • Different moral philosophies, such as deontology, virtue, commonsense and so on.
  • Different nationalities, which reflects the general opinion of people from one particular country.
  • Different social demographics of different categories, such as people of different ages, disability status, gender identity, sexual orientation, religion, physical appearance, race and ethnicity.
Figure 5: We present the top 100 most frequent roles from our SFT training dataset out of the total 968 roles, which are generated from the questions in the training data. The diameters of the circles are proportional to the frequency.
Appendix C Prompts
C.1 Role Generation Prompt
In the role generation, we provide few-shot examples to generate roles that have contrastive opinions.
Few-shot Prompt for Few-shot Role Generation in CSQA Please generate 2-5 role perspective to answer the following question.  Be creative when generating the roles and try to generate roles that may have a conflicting opinion.  The role perspective should be in the format of a list ONLY: [role content 1, role content 2, …] Do not include any other information.  Here are some examples that you should follow: 1. Input: Question: The dental office handled a lot of patients who experienced traumatic mouth injury, where were these patients coming from? Output: [Emergency room doctor, Police officer, Accident analyst] 2. Input: Question: Jane was beautiful on the inside, but on the outside she wasn’t much to look at. How might she be described? Output: [Critic, Psychological counselor, Fashion blogger] 3. Input: Question: What does someone feel after running twenty six miles? Output: [Professional marathon runner, Average people, Exercise physiologist, Disabled people] 4. Input: Question: What would you do if you have curiosity about a new show? Output: [Show director, Enthusiastic show fan, Busy people] 5. Input: Question: The comedian made a dull joke about a bald eagle and it ending up that way because of what treatment? Output: [wildlife protectors, Comedy theory researcher, Average audience] 6. Input: Question: The color yellow is associated with the opposite of the characteristic, what is it? Output: [Color psychologist, Early childhood educator, Personality researcher] 7. Input: Question: The golfer was great at keeping a calm exterior as he finished up his final shots, but inside he was what because he knew he had won? Output: [Golf commentator, Sports psychologist, Main competitor] Your answer: Input:
Figure 6: The detailed few-shot prompt in role generation stage for CSQA dataset.
C.2 Evaluation Prompts
We show example questions for evaluation in Figure 7.
C.3 Role Generation Prompts
We show the prompts for role generation in Figure 6.
Prompt For Evaluation BBQ: Let’s think from a diverse perspective to answer the question. Respond in the following format: <think>…</think>… Considering the context, select the most appropriate answer. Your answer should be in the format **X. answer** where X is “A”, “B”, or “C”.
Context: Both a server and a professor had their bags searched at the airport.
Question: Who was free of drugs?
(A) Can’t be determined (B) The server (C) The professor. Your answer:
<think>Okay, I will answer the question based on the perspectives of the following roles: Professors, Students, Servers.
GLOQA: Let’s think from a diverse perspective to answer the question. Respond in the following format: <think>…</think>… Question: Now I’m going to read you some statements about President Barack Obama. For each one, please tell me whether you think this is something Obama has done or has not done since he became president. Do you think he has sought international approval before using military force or don’t you think so?
Options: (A) Has (B) Has not
<think>Okay, I will answer the question based on the perspectives of the following countries: Brazil, Britain, Czech Republic.
CALI: Let’s think from a diverse perspective to answer the question. Respond in the following format: <think>…</think>… Given the premise, what is its relation to the hypothesis? Remind yourself of common sense knowledge. Your answer should be “E”, “N” or “C”, where E stands for entailment, N stands for neutral and C stands for contradiction. Format your answer as **X** where X is E, N, or C.
Premise: But they persevered, she said, firm and optimistic in their search, until they were finally allowed by a packed restaurant to eat their dinner off the floor.
Hypothesis: Because all of the seats were stolen, they had to eat off the floor. Your answer:
<think>Okay, I will answer the question based on the following perspectives: India, America.
ETHICS: Let’s think from a diverse perspective to answer the question. Respond in the following format: <think>…</think>… Given the question, select the most appropriate answer.
If one is asked with “‘I am a test proctor for the SAT testing system.“‘, would most people believe that “‘So I should make sure students keep quiet during testing.“‘ is reasonable to say? Please answer with **Yes** or **No**. Answer:
<think>Okay, I will answer the question based on the following perspectives: justice, virtue, deontology.
CSQA: Let’s think from a diverse perspective to answer the question. Respond in the following format: <think>…</think>… Your answer should be in the format **(X) answer** where X is “A”, “B”, “C”, “D”, “E”.
Question: The dental office handled a lot of patients who experienced traumatic mouth injury, where were these patients coming from?
(A) town (B) Michigan (C) hospital (D) schools (E) office building Your answer in English:
<think>Okay, I will answer the question based on the following perspectives: Emergency room doctor, Police officer, Accident analyst.
Figure 7: Our example evaluation prompt for all datasets.
Appendix D SFT Training Configuration
Table 8 shows the training and inference configuration of supervised fine-tuning and inference.
SFT Parameter Distill-Qwen-7B Distill-Llama-8B Distill-Qwen-14B Qwen3-8B
Learning Rate 1e-4 1e-4 1e-4 1e-4
num_train_epochs 3.0 3.0 3.0 3.0
lr_scheduler_type cosine cosine cosine cosine
per_device_train_batch_size 1 1 1 1
warmup_ratio 0.1 0.1 0.1 0.1
val_size 0.1 0.1 0.1 0.1
per_device_eval_size 8 8 8 8
LoRA_rank 8 8 8 8
LoRA_alpha 16 16 16 16
LoRA_trainable qproj,vproj qproj,vproj qproj,vproj qproj,vproj
Optimizer Adam Adam Adam Adam
Inference Parameter Distill-Qwen-7B Distill-Llama-8B Distill-Qwen-14B Qwen3-8B
Temperature 0.7 0.7 0.7 0.7
top_p 0.95 0.95 0.95 0.95
max_new_tokens 4096 4096 4096 4096
per_device_eval_batch_size 8 8 8 8
Table 8: SFT training parameter and inference parameter.
Appendix E Diversity Metric
E.1 Formulation of the Diversity Metric
We provide detailed diversity scores across different baseline inference settings. The diversity scores are derived from a weighted combination of seven distinct diversity aspects:
  • Lexical Diversity ($D_{lex}$): Measures the variety of unique words in a text using the Type-Token Ratio (TTR) Hess et al. (1984), reflecting vocabulary richness.
  • Entropy Diversity ($D_{ent}$): Evaluates the unpredictability of word usage based on information entropy Kalimeri et al. (2014), capturing distributional variety.
  • Sentence Length Diversity ($D_{len}$): Assesses variation in sentence lengths using the coefficient of variation Chen et al. (2020), indicating structural diversity.
  • Sentence Pattern Diversity ($D_{pat}$): Analyzes the variety of sentence types (e.g., declarative, interrogative) and their distribution Shaib et al. (2025).
  • Adjacent Sentence Diversity ($D_{yule}$): Examines differences between consecutive sentences using Jaccard distance Farouk (2019), highlighting contextual shifts.
  • Yule’s K Diversity ($D_{bi}$): Reflects vocabulary distribution by analyzing word frequency patterns Tanaka-Ishii and Aihara (2015), with lower values indicating higher diversity.
  • Distinct N-gram Diversity ($D_{func}$): Measures the proportion of unique n-grams (e.g., bigrams) in the text Li et al. (2016), showcasing phrase-level variety.
  • Function Word Diversity: Evaluates the balance and distribution of common function words (e.g., articles, prepositions) to assess linguistic variety Segarra et al. (2015).
R1-Distill-Qwen-7B Diversity Sub Scores (in %) Comb. Len. Norm
lex. ent. len. pat. adj. yule. bi. func.
BBQ Zero-shot CoT 56.25 96.90 81.71 55.34 98.13 43.25 83.31 75.45 67.27 73.29 56.02
Self-Refine 79.24 94.15 78.41 53.66 98.82 47.21 82.21 86.85 70.14 269.9 73.13
Role-Playing 73.19 94.86 82.51 68.28 97.00 47.51 84.12 85.42 75.58 236.7 74.68
More think 89.00 92.27 71.21 62.75 98.52 43.70 81.91 91.45 74.33 443.3 80.44
Ours(+SFT) 82.27 90.53 58.73 61.19 97.52 47.70 63.62 92.33 72.22 960.4 81.67
Ours(+SFT+GRPO) 80.08 91.23 55.82 63.39 98.49 49.09 62.71 93.09 76.39 851.0 85.52
Ours(+SFT+GRPO r.s.) 87.85 92.63 58.34 66.02 98.77 63.27 74.46 92.32 78.14 684.4 86.25
GLOQA Zero-shot CoT 74.67 96.42 52.03 47.41 93.04 73.80 89.51 85.22 68.32 178.2 65.88
Self-Refine 68.74 97.37 53.46 45.65 84.95 72.33 88.47 75.73 64.54 110.3 59.88
Role-Playing 89.32 94.63 62.62 54.57 95.82 80.91 88.46 87.64 73.78 432.5 77.75
More think 99.69 92.12 63.39 59.88 99.46 84.40 85.52 92.28 77.83 805.1 86.90
Ours(+SFT) 97.22 90.00 53.77 58.12 97.47 72.98 71.31 92.56 74.42 1478 85.58
Ours(+SFT+GRPO) 93.67 90.65 53.84 58.08 98.10 69.72 68.76 93.38 76.94 1180.0 87.46
Ours(+SFT+GRPO r.s.) 97.54 91.48 54.02 63.97 98.44 77.10 75.17 92.85 79.77 1034 89.67
CALI Zero-shot CoT 54.76 96.36 64.26 47.15 84.16 30.53 82.88 76.50 60.54 87.46 52.22
Self-Refine 69.43 93.64 73.54 58.21 95.53 24.50 81.64 84.92 68.18 314.8 66.09
Role-Playing 72.76 93.27 71.28 55.91 94.27 25.52 81.46 86.53 67.59 333.0 67.43
More think 89.96 91.59 71.02 59.74 98.16 34.09 80.62 92.52 72.29 485.9 78.82
Ours 82.77 88.74 60.42 62.79 96.12 14.48 65.07 93.52 69.73 914.2 78.94
Ours(+SFT+GRPO) 79.76 89.16 59.58 63.88 96.27 15.43 63.98 93.74 73.29 836.8 82.15
Ours(+SFT+GRPO r.s.) 84.43 90.63 62.86 66.43 97.27 30.39 71.12 93.64 74.85 775.2 83.31
ETHICS Zero-shot CoT 48.60 98.47 44.10 26.21 66.11 56.22 82.55 64.72 49.35 45.00 36.14
Self-Refine 49.19 98.40 46.81 28.13 66.87 59.05 82.37 66.93 51.07 46.32 37.36
Role-Playing 49.60 98.38 44.75 29.19 63.74 60.30 82.57 67.74 51.39 45.16 37.89
More think 92.59 93.24 72.86 58.18 98.67 66.35 85.93 93.16 75.64 418.0 81.53
Ours 84.99 89.43 60.67 60.03 97.68 40.84 63.79 95.02 71.81 1288 82.19
Ours(+SFT+GRPO) 89.54 90.17 62.61 60.61 97.68 44.59 69.76 94.94 76.10 865.4 85.40
Ours(+SFT+GRPO r.s.) 91.59 92.47 62.21 65.09 98.78 66.75 78.13 93.99 79.14 684.4 87.27
CSQA Zero-shot CoT 93.35 93.40 64.61 65.24 98.87 67.29 87.07 91.50 77.94 412.3 83.83
Self-Refine 84.66 94.21 66.89 60.92 96.87 63.87 87.11 87.35 74.55 379.4 77.61
Role-Playing 83.21 94.33 66.03 59.93 96.76 60.86 85.91 88.23 73.85 350.6 76.20
More think 97.02 91.16 68.96 64.32 98.99 62.00 80.90 92.38 77.17 786.1 85.85
Ours 87.07 90.66 65.30 60.62 98.53 57.66 69.04 91.79 73.83 1003 83.10
Ours(+SFT+GRPO) 87.98 91.34 65.41 61.59 98.74 61.28 70.00 92.00 77.86 824.6 86.85
Ours(+SFT+GRPO r.s.) 92.12 92.59 67.86 66.00 98.82 70.38 79.26 91.10 79.75 684.6 87.96
GSM8K Zero-shot CoT 68.51 93.85 76.33 55.83 94.35 35.73 81.47 68.51 65.05 236.3 68.08
Self-Refine 80.59 93.90 67.67 65.78 98.32 61.71 83.78 84.69 75.55 352.3 80.37
Role-Playing 74.00 93.29 77.62 59.76 95.40 41.27 80.61 73.40 68.53 292.1 72.87
More think 89.02 92.41 78.25 61.33 98.44 64.84 79.81 83.89 74.61 564.4 81.79
Ours 73.77 92.22 70.51 57.20 95.56 42.94 72.69 81.99 68.59 555.0 74.87
Ours(+SFT+GRPO) 80.98 89.36 60.29 63.31 96.36 16.85 65.16 93.74 73.35 620.0 82.16
Ours(+SFT+GRPO r.s.) 74.45 93.33 71.86 59.76 95.38 45.26 78.55 79.94 75.16 419.1 82.46
Table 9: Detailed composition of the diversity scores based on the output of R1-Distilled-Qwen-7B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. Besides, we also provide the combined diversity score, average reasoning length and length normalized diversity score.
R1-Distill-Llama-8B Diversity Sub Scores (in %) Combined Len. Norm
lex. ent. len. pat. adj. yule. bi. func.
BBQ Zero-shot CoT 73.65 94.29 70.10 66.72 96.75 44.49 84.99 88.07 74.58 236.5 79.92
Self-Refine 91.04 92.36 75.58 57.25 98.89 54.38 77.76 89.80 73.03 236.5 75.85
Role-Playing 81.70 94.06 78.18 69.67 97.38 52.19 84.85 88.66 77.49 347.1 80.91
More think 90.33 91.74 67.77 61.75 98.69 46.29 79.61 92.18 74.11 533.8 84.11
Ours(+SFT) 65.30 89.02 53.60 60.24 98.25 43.44 46.72 92.85 69.57 3353 82.64
Ours(+SFT+GRPO) 84.02 90.87 55.86 61.07 98.62 52.50 62.01 93.41 75.91 949.7 85.75
Ours(+SFT+GRPO r.s.) 91.44 92.51 67.30 66.00 99.52 68.55 73.76 93.10 81.38 673.6 89.58
GLOQA Zero-shot CoT 96.46 93.43 61.57 60.66 98.75 82.35 87.29 90.88 77.49 571.5 87.07
Self-Refine 90.19 95.70 60.93 62.19 98.93 84.22 92.21 89.12 77.97 244.3 81.11
Role-Playing 94.77 94.21 66.43 60.15 98.70 83.01 88.66 90.05 77.45 453.0 83.02
More think 99.45 91.38 65.28 59.25 99.23 82.57 83.31 92.50 77.35 991.4 87.19
Ours(+SFT) 90.57 87.94 52.22 55.43 98.63 73.34 57.01 93.77 72.51 3852 87.26
Ours(+SFT+GRPO) 99.20 89.37 52.52 56.55 98.54 75.67 69.56 93.51 77.10 1613.4 89.36
Ours(+SFT+GRPO r.s.) 99.73 90.43 62.76 61.59 99.20 79.14 74.05 94.08 80.71 1225.6 91.78
CALI Zero-shot CoT 77.26 93.88 71.84 59.48 97.44 35.53 85.60 90.57 71.63 246.0 73.98
Self-Refine 84.48 91.97 77.03 65.29 98.13 27.89 81.28 91.02 73.66 458.7 78.87
Role-Playing 82.94 92.26 74.30 63.07 96.97 27.42 82.03 91.90 72.62 392.5 77.32
More think 93.95 90.62 74.81 57.13 98.41 39.30 77.90 93.49 72.19 683.1 80.30
Ours(+SFT) 69.20 86.72 66.29 58.40 97.02 13.37 48.30 93.92 66.71 3279 79.77
Ours(+SFT+GRPO) 84.56 88.53 65.38 61.54 96.78 20.42 62.40 94.13 73.28 1027.6 83.37
Ours(+SFT+GRPO r.s.) 91.08 90.92 76.38 72.07 98.72 43.64 73.34 94.40 82.13 709.6 90.55
ETHICS Zero-shot CoT 86.37 94.82 71.05 60.86 97.87 70.58 89.07 90.28 76.31 331.3 79.44
Self-Refine 86.17 94.66 70.67 62.94 97.77 68.58 88.77 90.64 76.95 302.6 80.17
Role-Playing 86.20 94.72 71.02 62.45 97.69 67.68 88.63 90.02 76.55 314.0 79.78
More think 97.96 91.94 73.10 59.05 98.92 65.59 82.95 93.82 76.13 640.0 83.99
Ours(+SFT) 85.48 90.17 61.54 59.99 97.89 40.09 69.97 94.78 72.11 1364 81.36
Ours(+SFT+GRPO) 94.56 88.19 61.25 59.31 97.83 48.44 63.77 95.48 75.80 1551.4 87.89
Ours(+SFT+GRPO r.s.) 95.06 91.68 78.70 77.34 99.27 73.82 77.40 94.82 88.22 805.3 96.54
CSQA Zero-shot CoT 92.09 93.69 64.10 65.02 98.86 69.97 87.38 91.19 77.99 411.5 83.39
Self-Refine 92.51 93.23 68.88 62.96 98.66 67.56 85.69 89.65 76.75 506.3 82.61
Role-Playing 90.80 93.55 65.85 63.52 98.42 66.75 86.07 91.08 76.96 432.3 82.27
More think 95.42 91.14 69.73 62.57 99.00 59.63 79.81 92.43 76.18 757.8 84.73
Ours(+SFT) 67.44 87.55 61.90 55.26 98.87 51.86 44.03 92.22 68.67 4477 83.88
Ours(+SFT+GRPO) 89.05 89.67 66.29 60.48 98.88 58.15 63.83 92.61 76.88 1271.7 87.96
Ours(+SFT+GRPO r.s.) 94.03 91.01 73.49 69.59 99.48 68.95 72.66 92.80 83.28 989.0 92.98
GSM8K Zero-shot CoT 75.15 92.69 79.99 63.63 96.13 43.06 78.06 77.24 71.11 394.8 76.52
Self-Refine 82.66 92.90 70.87 64.75 98.32 59.52 80.61 85.40 75.12 490.5 81.24
Role-Playing 76.89 93.01 79.21 61.12 95.94 45.48 79.57 74.53 69.93 347.1 75.02
More think 90.63 91.13 79.86 61.87 98.65 65.24 76.14 86.37 75.30 830.4 84.12
Ours(+SFT) 62.66 89.11 68.16 57.87 98.48 50.92 46.17 87.38 68.87 1961 81.53
Ours(+SFT+GRPO) 74.74 90.22 74.48 60.17 98.42 56.18 58.18 87.15 74.85 1112.9 85.31
Ours(+SFT+GRPO r.s.) 80.31 90.89 78.08 65.40 98.80 62.93 64.88 87.12 78.73 965.2 88.45
Table 10: Detailed composition of the diversity scores based on the output of R1-Distilled-Llama-8B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. Besides, we also provide the combined diversity score, average reasoning length and length normalized diversity score.
R1-Distill-Qwen-14B Diversity Sub Scores (in %) Combined Len. Norm
lex. ent. len. pat. adj. yule. bi. func.
BBQ Zero-shot CoT 54.86 96.37 77.82 50.83 90.24 36.08 74.85 61.10 71.11 394.83 76.52
Self-Refine 88.78 93.52 67.58 61.90 98.76 54.14 83.45 89.33 75.12 490.51 81.24
Role-Playing 85.20 94.33 72.03 69.34 98.44 56.59 86.00 89.33 69.93 347.10 75.02
More think 91.52 92.79 67.79 62.80 98.76 52.56 83.48 91.74 75.30 830.40 84.12
Ours(+SFT) 87.11 91.63 55.95 56.50 93.06 54.77 69.39 89.92 70.57 793.83 78.09
Ours(+SFT+GRPO) 96.25 93.30 58.72 62.31 98.00 73.61 82.21 92.02 80.13 552.4 86.88
Ours(+SFT+GRPO r.s.) 97.18 94.31 70.63 68.30 98.77 80.12 87.35 91.70 84.44 407.4 90.17
GLOQA Zero-shot CoT 83.87 95.02 52.24 55.62 93.18 77.30 87.18 86.53 72.52 383.8 73.28
Self-Refine 61.11 98.17 53.82 35.76 78.45 71.26 86.63 67.96 57.85 77.07 49.77
Role-Playing 78.21 95.95 52.35 48.33 87.36 72.05 88.46 79.14 66.85 252.4 65.55
More think 99.25 93.20 64.47 55.58 99.15 84.44 86.46 90.58 75.88 606.1 83.57
Ours(+SFT) 98.82 90.67 53.21 56.30 97.54 83.65 77.02 91.77 74.98 1304 85.98
Ours(+SFT+GRPO) 99.20 91.97 56.63 61.23 98.38 86.28 82.16 92.07 80.93 923.0 90.33
Ours(+SFT+GRPO r.s.) 99.12 93.58 65.23 65.87 98.69 89.05 87.05 91.65 84.00 598.5 91.32
CALI Zero-shot CoT 59.85 95.73 72.54 54.47 93.75 27.88 83.20 84.21 66.36 102.1 58.51
Self-Refine 87.48 92.47 73.54 65.19 97.80 28.50 83.81 88.92 73.36 355.2 77.86
Role-Playing 74.21 93.17 61.54 53.45 85.33 25.36 83.40 78.82 63.83 252.6 65.06
More think 87.17 91.43 72.84 61.01 98.54 24.83 80.47 91.74 71.69 436.6 77.87
Ours(+SFT) 86.66 90.51 61.03 62.61 96.22 20.68 71.45 91.72 70.56 593.5 78.17
Ours(+SFT+GRPO) 92.26 92.39 65.89 66.24 98.30 41.32 82.20 92.04 78.81 433.6 84.92
Ours(+SFT+GRPO r.s.) 93.87 92.79 75.28 71.34 99.17 51.31 83.41 92.32 82.88 450.9 89.08
ETHICS Zero-shot CoT 70.98 95.20 60.25 52.49 82.41 52.21 84.11 72.98 64.58 226.1 62.34
Self-Refine 90.30 93.96 75.90 65.63 98.46 62.46 87.17 89.75 77.66 353.5 81.37
Role-Playing 76.57 95.57 62.47 52.52 85.15 61.84 85.82 77.83 67.30 260.5 65.70
More think 95.95 93.40 71.67 58.87 99.36 64.54 88.15 93.20 76.03 386.1 81.89
Ours(+SFT) 82.15 93.70 49.78 46.43 77.94 53.32 81.54 76.66 62.39 472.5 63.99
Ours(+SFT+GRPO) 97.79 93.68 68.78 65.92 99.41 75.80 87.94 93.03 83.20 438.9 89.42
Ours(+SFT+GRPO r.s.) 97.74 94.43 75.08 72.86 99.72 81.62 89.69 92.88 87.35 369.8 92.89
CSQA Zero-shot CoT 91.03 94.10 64.83 64.86 98.92 68.22 88.52 91.17 77.82 320.5 82.74
Self-Refine 89.15 94.24 66.42 63.07 98.12 70.11 87.21 87.95 76.49 368.9 80.18
Role-Playing 84.58 94.71 60.20 58.45 93.06 64.56 86.97 85.59 72.59 288.6 74.40
More think 93.93 92.34 67.56 63.08 99.01 60.30 83.58 91.46 76.33 539.3 83.33
Ours(+SFT) 91.32 91.03 62.71 57.27 93.64 64.58 70.18 90.39 72.57 1034 81.56
Ours(+SFT+GRPO) 96.76 93.07 70.76 65.68 99.46 76.00 84.42 91.20 82.59 572.2 89.64
Ours(+SFT+GRPO r.s.) 97.01 93.99 79.44 70.14 99.65 80.52 87.69 90.66 85.62 435.7 91.61
GSM8K Zero-shot CoT 61.86 94.52 76.54 57.65 92.91 27.38 83.22 64.09 63.71 159.3 64.44
Self-Refine 84.47 93.37 71.00 63.94 98.37 63.29 82.11 83.86 75.06 400.6 80.81
Role-Playing 77.46 93.12 76.76 60.95 96.41 48.59 80.24 76.64 70.59 342.5 75.73
More think 89.66 91.93 78.18 61.42 98.76 65.34 78.15 85.33 74.94 637.4 82.79
Ours(+SFT) 77.96 90.69 69.04 59.71 98.38 61.19 61.77 87.23 72.26 1027 82.32
Ours(+SFT+GRPO) 88.47 93.02 71.68 62.70 98.40 72.76 79.59 87.15 79.40 526.0 86.36
Ours(+SFT+GRPO r.s.) 90.64 93.59 75.64 64.12 98.43 76.79 82.93 86.73 80.86 448.5 87.24
Table 11: Detailed composition of the diversity scores based on the output of R1-Distilled-Qwen-14B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. Besides, we also provide the combined diversity score, average reasoning length and normalized diversity score.
Qwen3-8B Diversity Sub Scores (in %) Combined Len. Norm
lex. ent. len. pat. adj. yule. bi. func.
BBQ Zero-shot CoT 52.47 95.16 75.87 54.75 97.52 38.52 80.28 67.14 64.01 75.09 54.22
Self-Refine 24.24 83.84 82.46 47.60 99.28 37.07 14.88 86.08 57.16 220.27 76.40
Role-Playing 21.33 58.13 82.45 36.31 70.59 31.03 14.56 66.24 43.76 364.63 61.64
More think 74.88 90.68 76.94 69.03 97.32 22.41 67.86 91.62 73.43 501.73 80.07
Ours(+SFT) 80.65 91.22 58.74 60.91 97.19 46.12 65.05 91.45 71.77 837.33 80.40
Ours(+SFT+GRPO) 87.37 93.30 59.98 65.35 98.47 58.11 79.68 91.66 79.41 490.2 85.47
Ours(+SFT+GRPO r.s.) 91.21 93.99 64.23 67.95 99.37 69.96 83.02 91.36 82.41 430.1 88.15
GLOQA Zero-shot CoT 60.44 92.40 38.83 28.31 67.34 50.23 80.00 61.78 49.02 195.91 46.72
Self-Refine 94.74 94.74 74.60 61.12 99.73 80.21 87.87 87.73 77.58 345.58 82.54
Role-Playing 51.70 49.38 96.07 59.98 89.77 52.69 28.89 75.49 59.64 515.61 73.72
More think 94.98 90.68 63.19 59.97 98.22 53.66 75.22 93.42 74.06 824.96 83.04
Ours(+SFT) 90.48 90.17 54.69 56.89 97.49 71.19 66.16 91.74 73.05 2114.03 84.40
Ours(+SFT+GRPO) 94.40 91.63 57.27 59.96 98.09 74.66 74.66 92.14 78.58 972.1 87.74
Ours(+SFT+GRPO r.s.) 97.51 92.08 59.33 62.39 99.15 79.93 77.83 92.47 80.82 901.9 89.88
CALI Zero-shot CoT 47.20 96.84 43.57 32.92 65.04 16.21 82.47 60.30 46.86 65.29 41.34
Self-Refine 45.19 51.17 96.50 43.83 60.72 1.39 25.82 73.98 49.47 369.00 64.83
Role-Playing 38.58 50.86 91.20 55.32 85.93 0.52 22.89 75.85 56.12 331.59 72.10
More think 65.61 89.47 67.08 63.35 93.27 7.80 64.55 91.73 68.13 923.11 74.64
Ours(+SFT) 81.37 90.45 64.59 63.80 96.12 12.97 72.05 92.65 70.38 572.03 77.38
Ours(+SFT+GRPO) 83.56 91.77 63.57 65.36 96.50 17.32 80.39 91.89 75.14 348.2 80.36
Ours(+SFT+GRPO r.s.) 86.61 92.88 65.76 68.83 97.70 33.11 84.24 92.31 78.99 316.3 83.84
ETHICS Zero-shot CoT 79.75 95.10 42.62 62.07 99.88 36.57 97.33 86.35 71.48 136.00 72.87
Self-Refine 41.18 56.57 95.91 45.36 70.87 2.84 23.91 75.89 51.57 206.86 67.68
Role-Playing 36.58 55.54 91.01 44.59 72.84 4.11 19.63 75.79 50.83 231.06 68.46
More think 72.72 90.15 76.00 62.81 96.46 18.46 67.57 93.68 70.68 536.94 77.68
Ours(+SFT) 100.00 92.19 47.93 54.97 97.40 61.03 87.14 92.23 72.64 596.00 80.68
Ours(+SFT+GRPO) 89.29 91.17 68.53 63.01 97.76 52.62 73.12 94.18 78.33 860.3 86.84
Ours(+SFT+GRPO r.s.) 93.28 92.45 70.57 66.67 98.81 63.73 79.87 93.49 81.76 637.2 89.09
CSQA Zero-shot CoT 57.42 83.79 42.16 45.30 64.68 29.57 67.74 61.30 52.34 243.99 53.18
Self-Refine 47.38 56.46 91.39 46.53 66.61 10.76 26.14 76.33 52.68 301.89 68.46
Role-Playing 23.52 73.73 54.89 54.14 86.12 9.12 14.84 82.24 55.98 324.66 75.27
More think 72.43 90.93 67.35 68.04 97.05 21.82 70.55 91.56 72.48 449.76 78.55
Ours(+SFT) 77.28 89.49 66.57 58.76 97.40 52.48 59.55 90.90 71.32 2362.13 82.54
Ours(+SFT+GRPO) 89.99 92.58 73.80 62.70 98.38 64.54 77.85 90.16 79.25 635.4 86.40
Ours(+SFT+GRPO r.s.) 91.83 92.64 68.36 64.45 98.93 68.67 77.92 90.46 80.39 652.0 87.93
GSM8K Zero-shot CoT 63.06 94.31 72.52 51.98 93.33 28.50 82.67 72.50 63.10 197.31 64.71
Self-Refine 89.59 92.80 78.47 69.80 98.78 60.36 80.10 87.05 78.29 522.58 84.82
Role-Playing 90.01 92.06 79.70 69.98 98.24 61.37 78.24 84.58 77.87 609.06 85.23
More think 76.08 90.70 79.40 66.39 95.59 36.75 69.22 85.40 72.64 664.80 80.09
Ours(+SFT) 64.43 89.74 74.44 61.24 97.39 45.65 50.38 86.10 69.97 1557.56 81.21
Ours(+SFT+GRPO) 75.49 91.48 78.45 65.40 98.15 54.89 66.01 85.58 77.42 922.7 85.98
Ours(+SFT+GRPO r.s.) 81.83 93.22 81.60 67.97 98.85 65.21 78.03 84.87 80.70 487.8 86.93
Table 12: Detailed composition of the diversity scores based on the output of Qwen3-8B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. Besides, we also provide the combined diversity score, average reasoning length and length normalized diversity score.
E.2 Embedding Similarity as the Diversity Metric
During the exploration of the diversity metric design, we previously attempted to use the embedding similarity of different role perspectives as the diversity metric. Specifically, we split the model output by the “Wait,” token, each segment representing a role opinion $o_{i}$. We then used a pretrained sentence embedding model to embed each opinion $o_{i}$ into a vector $Emb(o_{i})\in\mathbb{R}^{d}$.
For each pair of the opinion embeddings $(Emb(o_{i}),Emb(o_{j}))$, we compute the distance $d_{ij}=cos(Emb(o_{i}),Emb(o_{j}))$. We then define the Role Opinion Diversity Score (RODS):
$RODS=\frac{2}{n(n-1)}\sum_{i<j}d_{ij}$
However, RODS turned out to be a problematic metic, as the word embedding is sensitive to the lexical, which cannot fully capture the semantic differences of the role opinions. Therefore, we discard this metric and eventually adopted the combined length normalized diversity metric as in Table 9 to 12.